Lucido Products markets two computer games: Claimjumper and Makeover. A contribution format income statement for a recent month for the two games appears below:

Claimjumper Makeover Total
Sales $ 110,000 $ 55,000 $ 165,000
Variable expenses 27,400 5,600 33,000
Contribution margin $ 82,600 $ 49,400 132,000
Fixed expenses 93,360
Net operating income $ 38,640

Required:

1. What is the overall contribution margin (CM) ratio for the company?

2. What is the company's overall break-even point in dollar sales?

3. Prepare a contribution format income statement at the company's break-even point that shows the appropriate levels of sales for the two products.
Answer #1
1
Total contribution margin 132000
Divide by Total sales 165000
Overall Contribution margin (CM) ratio 80%
2
Fixed expenses 93360
Contribution margin (CM) ratio 80%
Overall break-even point 116700
3
Lucido Products
Contribution Income Statement
Claimjumper Makeover Total
Sales 77800 38900 116700
Variable expenses 19379 3961 23340
Contribution margin 58421 34939 93360
Fixed expenses 93360
Net operating income 0
Workings:
Sales mix:
Claimjumper = 110000/165000 = 2/3
Makeover = 55000/165000 = 1/3
Sales at break even:
Claimjumper = 116700*2/3= 77800
Makeover = 116700*1/3= 38900
Variable expenses at break even:
Claimjumper = 27400/110000*77800 = 19379
Makeover = 5600/55000*38900 = 3961